# AGZO: Activation-Guided Zeroth-Order Optimization for LLM Fine-Tuning

### v1

@@ 1 Introduction
Despite this progress, existing ZO fine-tuning baselines all follow a black-box perturbation paradigm: perturbations are sampled from data-independent Gaussian distributions defined solely by parameter dimensions. MeZO uses isotropic full-space perturbations (Malladi et al., 2023), and recent variant LOZO explores low-rank perturbations motivated by spectral properties of gradients (Chen et al., 2025b). Across these approaches, the perturbation distribution is typically independent of the internal representations produced by the current forward pass. This is a missed opportunity, as the forward evaluation exposes structural information that is intrinsically linked to the true gradient direction.
@@
This paper is motivated by a simple observation about backpropagation in linear layers of LLMs: the weight gradient on a mini-batch is completely determined by the upstream signals flowing into the layer and the activations produced in the forward pass. As a consequence, the gradient directions of a linear layer are confined to the subspace spanned by the mini-batch activations, rather than being arbitrary in the full parameter space. In parallel, a growing body of empirical and theoretical work suggests that adaptation during fine-tuning is effectively low-dimensional and often admits low-rank structure, which has been exploited by parameter-efficient and low-rank update methods (Aghajanyan et al., 2021; Li et al., 2018; Hu et al., 2022; Hao et al., 2024). These insights motivate a simple principle for ZO fine-tuning: instead of perturbing weights in unconstrained random directions, one should concentrate perturbations within an activation-informed low-dimensional subspace revealed by the forward pass.
Guided by this principle, we propose Activation-Guided Zeroth-Order optimization (AGZO). Each iteration of AGZO starts with the standard forward pass, which both evaluates the loss and produces the mini-batch activations. For each linear layer, AGZO then computes a compact set of dominant activation directions using a lightweight power iteration process (Golub & Van Loan, 2013; Miyato et al., 2018). AGZO samples low-rank, activation-guided perturbations constrained to this subspace, and uses them to form the usual forward-only ZO update. For nonlinear trainable layers, AGZO falls back to standard Gaussian perturbations to preserve general applicability across architectures. Crucially, AGZO does not retain activations across iterations: activations are discarded immediately after subspace extraction, and only compact subspace information is kept, thereby maintaining the memory advantages of ZO fine-tuning. Compared with existing perturbation baselines, AGZO exploits per-iteration activation structure to construct more informative perturbations, improving update quality while retaining the same memory efficient character.
@@
We provide a theoretical framework explaining the efficacy of activation guidance. By analyzing the alignment between the estimated and true gradients, we prove that AGZO generates update directions that are provably closer to the true gradient than those produced by random full-space perturbation baselines. We interpret AGZO as optimizing a subspace-smoothed objective and analyze how well its estimator aligns with the true gradient in expectation. Under an activation spectral concentration condition—formalizing that most gradient energy lies in the leading singular directions of the activation matrix (Gur-Ari et al., 2018; Papyan, 2020)—we show that AGZO achieves a larger expected cosine similarity to the true gradient than full-space random perturbation baseline. This result formalizes the intuition that activation-informed subspaces focus the optimization on directions carrying meaningful gradient signals, thereby enhancing the effectiveness of each update step
@@
  • We identify and formalize a fundamental structural link between gradients and activations in linear layers and empirically demonstrate that, during LLM fine-tuning, the gradient signal concentrates in a low-dimensional subspace revealed by the forward-pass activations.
  • Building on this insight, we propose AGZO, a zeroth-order fine-tuning method that extracts compact activation subspaces on the fly and uses them to construct low-rank, activation-guided perturbations.
  • We develop a principled theoretical framework that explains why activation guidance improves ZO optimization. We show that AGZO can be viewed as optimizing a subspace-smoothed objective, and its update directions are provably more aligned with the true gradient than random perturbation methods under activation spectral concentration.
  • We conduct experiments on Qwen3 models that jointly demonstrate (i) improved end-to-end fine-tuning performance over prior ZO baselines, (ii) consistently stronger gradient alignment with the true backpropagation direction, and (iii) a peak GPU memory footprint that remains essentially unchanged relative to standard forward-only ZO methods across varying batch size and sequence length.
2 Background: Zeroth-Order Fine-Tuning Baselines
@@
$\min_{Wa\in\mathbb{R}^{d}}F(W)\;\triangleq\;\mathbb{E}_{B\sim\mathcal{D}}\bigl[f(W;B)\bigr],$ (1)
where $W$ denotes the parameters of a LLM, $\mathcal{D}$ is a data distribution over minibatches $B$, and $f(W;B)$ is the empirical loss. First-order methods estimate $\nabla F(W)$ via backpropagation, which requires storing forward activations and thus dominates the training memory footprint. Zeroth-order methods approximate gradient directions using only function evaluations and avoid backpropagation, making them attractive for memory-limited fine-tuning of large models (Zhang et al., 2024; Malladi et al., 2023). In this section we briefly review two representative ZO baselines for LLM fine-tuning: MeZO and LOZO (Malladi et al., 2023; Chen et al., 2025b).
2.1 MeZO: Memory-Efficient Full-Space Perturbations
MeZO adapts classical Gaussian-smoothing ZO estimators to the LLM fine-tuning setting. Let $u\sim\mathcal{N}(0,I_{d})$ be a standard Gaussian perturbation, organized layer-wise as $u=(U_{1},\dots,U_{L})$ where each $U_{\ell}$ has the same shape as $W_{\ell}$. Given a smoothing parameter $\mu>0$ and a minibatch $B$, MeZO performs two forward passes to evaluate $f(W+\mu u;B)$ and $f(W;B)$ and constructs the finite difference estimator
@@ 2.1 MeZO: Memory-Efficient Full-Space Perturbations
This estimator can be interpreted as the gradient of a Gaussian-smoothed objective $F_{\mu}(W)=\mathbb{E}_{u}[F(W+\mu u)]$ and is used as a surrogate gradient in a standard optimizer.
To remain memory efficient, MeZO never stores the full perturbation tensor $u$ explicitly: it perturbs parameters in place and records only the random seed needed to regenerate $u$ when forming Eq (2). This design reduces fine-tuning memory by roughly a factor of four relative to first order methods while maintaining competitive performance on downstream tasks (Zhang et al., 2024; Gautam et al., 2024). However, $u$ is supported on the entire parameter space, and empirical studies suggest that layer-wise gradients in LLMs are effectively low-rank (Aghajanyan et al., 2021; Li et al., 2018; Chen et al., 2025b). Full-space isotropic perturbations may spend a substantial portion of the query budget exploring directions that carry little gradient energy.
2.2 LOZO: Low-Rank Zeroth-Order Estimation
LOZO aims to better match the observed low-rank structure of gradients by introducing a matrix-wise low-rank ZO estimator (Chen et al., 2025b). For each layer $\ell$, LOZO samples random Gaussian factors $U_{\ell}\in\mathbb{R}^{d_{out}\times r_{\ell}}$ and $V_{\ell}\in\mathbb{R}^{d_{in}\times r_{\ell}}$ with $r_{\ell}\ll\min\{d_{out},d_{in}\}$, and forms the rank-$r_{\ell}$ perturbation
@@ 2.2 LOZO: Low-Rank Zeroth-Order Estimation
so that the full perturbation is $\Delta=(\Delta_{1},\dots,\Delta_{L})$. LOZO defines the low-rank gradient estimator (LGE) as
@@
where $r=\{r_{\ell}\}_{\ell=1}^{L}$ and the division is understood layer-wise as $\Delta_{\ell}/r_{\ell}$. For each layer $\ell$, the estimate $\widehat{g}^{\text{LOZO}}_{\mu}(W;B)_{\ell}$ has rank at most $r_{\ell}$. Compared with MeZO, LOZO enforces a low-rank structure that more closely resembles FO gradients in LLM fine-tuning. At the same time, the additional memory cost is small: only the factors $U_{\ell}$ and $V_{\ell}$ (or their random seeds) need to be stored to reconstruct $\Delta_{\ell}$. This makes LOZO a natural low-rank ZO baseline for large-scale fine-tuning.
@@ 2.3 Limitations of Existing ZO Baselines
Figure 1: Structural analysis of gradients and activations. (a) & (b) Singular value spectra of gradients and activations. (c) Cosine similarity between the true gradient and its projection onto the activation subspace.
3 Gradient and Activation Structure as Motivation
This section analyzes the structural properties of gradients in linear layers, revealing both determining links to forward activations and inherent low-rank characteristics. We show that (i) gradients of linear layers admit a simple matrix factorization involving the activation matrices, (ii) both gradients and activations exhibit strong low-rank behavior, and (iii) the gradient row-space is almost entirely contained in the activation column space. These observations motivate constructing zeroth-order perturbations inside an activation-informed low-rank subspace rather than in the full parameter space.
3.1 Notation and Layer-Wise Gradient Forms
We focus on linear layers in LLMs, such as the projection matrices within self-attention mechanisms and fully connected layers in feed-forward networks. These layers dominate the model scale, accounting for most of the trainable parameters in many architectures (Kaplan et al., 2020; Vaswani et al., 2017; Yang et al., 2025). Consider a network with $L$ layers and linear parameters $W_{1},\dots,W_{L}$.
In Transformer-based architectures, the input consists of a minibatch of sequences. Let $b$ denote the batch size and $T$ the sequence length. For a linear layer $\ell$ with weight matrix $W_{\ell}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, the layer operates position-wise on the token representations. Let $h_{\ell,t}^{(i)}\in\mathbb{R}^{d_{\text{in}}}$ denote the input activation for the $t$-th token of the $i$-th sequence ($1\leq i\leq b,1\leq t\leq T$). The pre-activation output $z_{\ell,t}^{(i)}$ is given by:
$z_{\ell,t}^{(i)}\;=\;W_{\ell}h_{\ell,t}^{(i)}.$ (5)
Let $q_{\ell,t}^{(i)}=\partial f/\partial z_{\ell,t}^{(i)}$ denote the upstream gradient with respect to the pre-activation for that specific token. By the chain rule, the gradient of the loss with respect to $W_{\ell}$ is the sum of contributions from all tokens in the minibatch:
$\nabla_{W_{\ell}}f(W;B)\;=\;\sum_{i=1}^{b}\sum_{t=1}^{T}q_{\ell,t}^{(i)}\bigl(h_{\ell,t}^{(i)}\bigr)^{\top}.$ (6)
To express this in compact matrix form, we conceptually flatten the batch and sequence dimensions into a single dimension $m=b\times T$, representing the total number of tokens in the minibatch. We define the aggregated activation matrix $H_{\ell}\in\mathbb{R}^{d_{\text{in}}\times m}$ and gradient matrix $Q_{\ell}\in\mathbb{R}^{d_{\text{out}}\times m}$ by concatenating the column vectors for all tokens:
$H_{\ell}=\bigl[h_{\ell,1}^{(1)},\dots,h_{\ell,T}^{(b)}\bigr],Q_{\ell}=\bigl[q_{\ell,1}^{(1)},\dots,q_{\ell,T}^{(b)}\bigr],$ (7)
so that the minibatch gradient takes the matrix form
$\nabla_{W_{\ell}}f(W;B)\;=\;Q_{\ell}H_{\ell}^{\top},$ (8)
which implies that the row-space of the layer-wise gradient is contained in the column space of the corresponding activation matrix:
@@ 3.1 Notation and Layer-Wise Gradient Forms
To quantify how tightly the gradient concentrates in this activation subspace, Fig. 1(c) plots a bar chart of the cosine similarity between the true gradient and its orthogonal projection onto the subspace spanned by the forward activations. Results are obtained when fine-tuning GPT-2 (Radford et al., 2019) on the SST-2 dataset (Wang et al., 2018). Concretely, we perform an SVD of the activation matrix $H_{\ell}$ and use the leading $r$ singular vectors to define a rank-$r$ activation subspace; we then project the gradient onto this subspace and compute the cosine similarity between the original gradient and the projected one. We report results for $r=1,10$ and for the full activation subspace ($r=750$). Across layers, the cosine similarity is typically close to 1 when $r\geq 10$, indicating that almost all gradient energy lies in the subspace spanned by the forward activations.
@@ 3.2 Low-Rank Structure of Gradients and Activations
The matrix factorization in Eq. (8) reveals the structural dependency of gradients on forward activations. While the weight matrices $W_{\ell}$ typically reside in high-dimensional spaces, empirical observations suggest that the actual information content typically concentrates in a much smaller subspace.
To examine this structure more concretely, we compute the singular values of $\nabla_{W_{\ell}}f(W;B)$ via SVD and plot them (on a log scale) for a few representative layers and training steps. Figure 1(a) shows that the singular values decay rapidly: the spectrum is far from flat, and a small number of leading singular directions dominate the rest. This supports the view that layer-wise gradients are effectively low-rank.
We observe a similar spectral phenomenon for the forward activation matrices. For each $H_{\ell}$, we compute its singular values and visualize them in Figure 1(b). Again, the singular values exhibit pronounced decay, indicating that the majority of the activation energy is concentrated along a few dominant directions.
Together with the overlap ratios $\rho_{\ell}$ in Figure 1(c), these results show that layer-wise gradient information is concentrated in a low-dimensional subspace that is almost entirely determined by the corresponding activation matrix, and that this activation subspace itself has a rapidly decaying spectrum and can be captured by a small number of leading directions. From a zeroth-order perspective, it is therefore natural to restrict perturbations to a low-rank subspace extracted from forward activations, rather than sampling arbitrary directions in the full parameter space. In the next section we instantiate this idea in an activation-guided ZO method that constructs perturbations inside such activation-informed subspaces.
@@ 4 Activation-Guided Zeroth-Order Optimization
We now introduce Activation-Guided Zeroth-Order optimization (AGZO). For linear layers, AGZO perturbs weights inside an activation-guided low-rank subspace; for nonlinear layers, AGZO simply uses full Gaussian perturbations as in MeZO. Each iteration reuses the standard forward pass to both evaluate the loss and extract dominant activation directions, without storing activation matrices across iterations.

### v2

@@ 1 Introduction
Despite this progress, existing ZO fine-tuning baselines all follow a black-box perturbation paradigm: perturbations are sampled from data-independent Gaussian distributions defined solely by parameter dimensions. MeZO uses isotropic full-space perturbations (Malladi et al., 2023), and recent variant LOZO explores low-rank perturbations motivated by spectral properties of gradients (Chen et al., 2025b). Across these approaches, the perturbation distribution is typically independent of the internal representations produced by the current forward pass, which overlooks the structural information exposed by the forward evaluation that is intrinsically linked to the true gradient direction.
@@
Motivated by a simple observation that the weight gradient on a mini-batch is determined by the upstream signals flowing into the layer and the activations produced in the forward pass, we demonstrate that the gradient directions of a linear layer are confined to the subspace spanned by the mini-batch activations, rather than being arbitrary in the full parameter space. Moreover, based on the empirical and theoretical insights that adaptation during fine-tuning is effectively low-dimensional and often admits low-rank structure  (Aghajanyan et al., 2021; Li et al., 2018; Hu et al., 2022; Hao et al., 2024), we propose a simple principle for ZO fine-tuning: instead of perturbing weights in unconstrained random directions, one should concentrate perturbations within an activation-informed low-dimensional subspace revealed by the forward pass.
Guided by this principle, we propose Activation-Guided Zeroth-Order optimization (AGZO). AGZO samples low-rank, activation-guided perturbations constrained to the subspace, and uses them to form the ZO update. For nonlinear trainable layers, AGZO falls back to standard Gaussian perturbations to preserve general applicability across architectures.
Moreover, AGZO utilizes several strategies to reduce the memory consumption. First, we conduct a lightweight power iteration process in subspace construction (Golub & Van Loan, 2013; Miyato et al., 2018). Second, AGZO keeps compact subspace information and releases in-memory cached activation values immediately after subspace extraction, thereby maintaining the memory advantages of ZO fine-tuning. Compared with existing perturbation baselines, AGZO exploits per-iteration activation structure to construct more informative perturbations, improving update quality while retaining the same memory-efficient character.
@@
We theoretically evaluate the efficacy of our proposed activation-guidance methodology. By analyzing the alignment between the estimated and true gradients, given that most gradient energy lies in the leading singular directions of the activation matrix (Gur-Ari et al., 2018; Papyan, 2020), we prove that AGZO achieves a larger expected cosine similarity to the true gradient than the full-space random perturbation baselines. This result formalizes the intuition that activation-informed subspaces focus the optimization on directions carrying meaningful gradient signals, thereby enhancing the effectiveness of each update step.
@@
  • We propose AGZO, a zeroth-order fine-tuning method that extracts compact activation subspaces on the fly and uses them to construct low-rank, activation-guided perturbations. We identify and formalize a fundamental structural link between gradients and activations in linear layers and empirically demonstrate that, during LLM fine-tuning, the gradient signal concentrates in a low-dimensional subspace revealed by the forward-pass activations.
  • We theoretically demonstrate that our proposed activation guidance method improves ZO optimization. We show that AGZO can be viewed as optimizing a subspace-smoothed objective, and its update directions are provably more aligned with the true gradient than random perturbation methods under activation spectral concentration.
  • We conduct experiments on Qwen3 models that jointly demonstrate (i) consistently stronger gradient alignment with the true backpropagation direction, (ii) improved end-to-end fine-tuning performance over prior ZO baselines, and (iii) a peak GPU memory footprint that remains essentially unchanged relative to standard forward-only ZO methods across varying batch size and sequence length.
2 Background: Zeroth-Order Fine-Tuning
@@
$\min_{W\in\mathbb{R}^{d}}F(W)\;\triangleq\;\mathbb{E}_{B\sim\mathcal{D}}\bigl[f(W;B)\bigr],$ (1)
where $W$ denotes the parameters of a LLM, $\mathcal{D}$ is a data distribution over minibatches $B$, and $f(W;B)$ is the empirical loss. First-order methods estimate $\nabla F(W)$ via backpropagation, which requires storing forward activations and thus dominates the training memory footprint (Zhang et al., 2024). Zeroth-order methods approximate gradient directions using only function evaluations and avoid backpropagation, making them attractive for memory-limited fine-tuning of large models (Zhang et al., 2024; Malladi et al., 2023). In this section we briefly review two representative ZO baselines for LLM fine-tuning: MeZO and LOZO (Malladi et al., 2023; Chen et al., 2025b).
2.1 Memory-Efficient Full-Space Perturbations
MeZO adapts classical Gaussian-smoothing ZO estimators to the LLM fine-tuning setting. Let $u\sim\mathcal{N}(0,I_{d})$ be a standard Gaussian perturbation, organized layer-wise as $u=(U_{1},\dots,U_{L})$ where each $U_{\ell}\in\mathbb{R}^{\frac{d}{L}}$ are layer-wise perturbation parameters. Given a small smoothing parameter $\mu>0$ and a minibatch $B$, MeZO performs two forward passes to evaluate $f(W+\mu u;B)$ and $f(W;B)$ and constructs the finite difference estimator
@@ 2.1 Memory-Efficient Full-Space Perturbations
This estimator can be interpreted as the gradient of a Gaussian-smoothed objective $F_{\mu}(W)=\mathbb{E}_{u}[F(W+\mu u)]$ and is used as a surrogate gradient in a standard optimizer (Nesterov & Spokoiny, 2017).
To remain memory efficient, MeZO never stores the full perturbation tensor $u$ explicitly: it perturbs parameters in place and records only the random seed needed to regenerate $u$ when forming Eq (2). This design reduces fine-tuning memory by roughly a factor of four relative to first-order methods while maintaining competitive performance on downstream tasks (Zhang et al., 2024; Gautam et al., 2024). However, $u$ is supported on the entire parameter space, and empirical studies suggest that layer-wise gradients in LLMs are effectively low-rank (Aghajanyan et al., 2021; Li et al., 2018; Chen et al., 2025b). Full-space isotropic perturbations may spend a substantial portion of the query budget exploring directions that carry little gradient energy.
2.2 Low-Rank Zeroth-Order Perturbations
LOZO aims to better match the observed low-rank structure of gradients by introducing a low-rank ZO estimator (Chen et al., 2025b). For each layer $\ell$, LOZO samples random Gaussian factors $U_{\ell}\in\mathbb{R}^{d_{\text{out}}\times r_{\ell}}$ and $V_{\ell}\in\mathbb{R}^{d_{\text{in}}\times r_{\ell}}$ with $r_{\ell}\ll\min\{d_{\text{out}},d_{\text{in}}\}$, and forms the rank-$r_{\ell}$ perturbation
@@ 2.2 Low-Rank Zeroth-Order Perturbations
so that the full perturbation is $\Delta=(\Delta_{1},\dots,\Delta_{L})$. LOZO defines the low-rank gradient estimator as
@@
where $r=\{r_{\ell}\}_{\ell=1}^{L}$ and the division is understood layer-wise as $\Delta_{\ell}/r_{\ell}$. Compared with MeZO, LOZO enforces a low-rank structure that more closely resembles FO gradients in LLM fine-tuning.
@@ 2.3 Limitations of Existing ZO Baselines
Figure 1: Structural analysis of gradients and activations. (a) Cosine similarity between the true gradient and its projection onto the activation subspace. (b) & (c) Singular value spectra of gradients and activations.
3 Structural Analysis of Gradients and Activations
This section analyzes the structural properties of gradients in linear layers, revealing both deterministic links to forward activations and inherent low-rank characteristics. We show that (i) gradients of linear layers admit a simple matrix factorization involving the activation matrices, (ii) both gradients and activations exhibit strong low-rank behavior, and (iii) the gradient row-space is almost entirely contained in the activation column space. These observations motivate constructing zeroth-order perturbations inside an activation-informed low-rank subspace rather than in the full parameter space.
3.1 Gradient Confinement in Activation Subspaces
We focus on linear layers in LLMs, such as the projection matrices within self-attention mechanisms and fully connected layers in feed-forward networks. These layers dominate the model scale, accounting for most of the trainable parameters in many architectures (Kaplan et al., 2020; Vaswani et al., 2017; Yang et al., 2025).
In Transformer-based architectures, the training data consists of minibatches of sequences. Let $b$ denote the batch size and $T$ the sequence length. While the model processes data as sequences, we can flatten the batch and sequence dimensions into a single dimension $m=b\times T$, representing the total number of tokens in the minibatch. For a linear layer $\ell$ with weight matrix $W_{\ell}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, we define the aggregated input activation matrix $H_{\ell}\in\mathbb{R}^{d_{\text{in}}\times m}$ by concatenating the activation vectors of all $m$ tokens. Similarly, let $Q_{\ell}\in\mathbb{R}^{d_{\text{out}}\times m}$ denote the matrix of upstream gradients with respect to the layer’s pre-activations.
Standard backpropagation computes the gradient of the loss with respect to $W_{\ell}$ by aggregating contributions across all tokens, which takes the compact matrix form:
$\nabla_{W_{\ell}}f(W;B)\;=\;Q_{\ell}H_{\ell}^{\top}.$ (5)
This factorization reveals a fundamental geometric property: the gradient matrix is formed by a linear combination of the columns of $H_{\ell}$, i.e., the row-space of the layer-wise gradient is strictly contained in the subspace spanned by the input activations:
@@ 3.1 Gradient Confinement in Activation Subspaces
We further quantify how tightly the gradient concentrates in this activation subspace. Fig. 1(a) plots a bar chart of the cosine similarity between the true gradient and its orthogonal projection onto the subspace spanned by the forward activations. Results are obtained when fine-tuning GPT-2 (Radford et al., 2019) on the SST-2 dataset (Wang et al., 2018). Concretely, we perform an SVD of the activation matrix $H_{\ell}$ and use the leading $r$ singular vectors to define a rank-$r$ activation subspace; we then project the gradient onto this subspace and compute the cosine similarity between the original gradient and the projected one. We report results for $r=1,10$ and for the full activation subspace ($r=750$). Across layers, the cosine similarity is typically close to 1 when $r\geq 10$, indicating that almost all gradient energy lies in the subspace spanned by the forward activations.
@@ 3.2 Low-Rank Structure of Gradients and Activations
The matrix factorization in Eq. (5) reveals the structural dependency of gradients on forward activations. While the weight matrices $W_{\ell}$ reside in high-dimensional spaces, empirical observations suggest that the actual information content typically concentrates in a much smaller subspace.
To examine this structure more concretely, we compute the singular values of $\nabla_{W_{\ell}}f(W;B)$ via SVD and plot them (on a log scale) for a few representative layers and training steps. Fig. 1(b) shows that the singular values decay rapidly: the spectrum is far from flat, and a small number of leading singular directions dominate the rest. This supports the view that layer-wise gradients are effectively low-rank.
We observe a similar spectral phenomenon for the forward activation matrices. For each $H_{\ell}$, we compute its singular values and visualize them in Fig. 1(c). Again, the singular values exhibit pronounced decay, indicating that the majority of the activation energy is concentrated along a few dominant directions.
These results show that layer-wise gradient information is concentrated in a low-dimensional subspace that is almost entirely determined by the corresponding activation matrix, and that this activation subspace itself has a rapidly decaying spectrum and can be captured by a small number of leading directions. From a zeroth-order perspective, it is therefore natural to restrict perturbations to a low-rank subspace extracted from forward activations, rather than sampling arbitrary directions in the full parameter space. In the next section we instantiate this idea in an activation-guided ZO method that constructs perturbations inside such activation-informed subspaces.
@@ 4 Activation-Guided Zeroth-Order Optimization
We now introduce Activation-Guided Zeroth-Order optimization (AGZO). For linear layers, AGZO perturbs weights inside an activation-guided low-rank subspace; for nonlinear layers, AGZO simply uses Gaussian perturbations. Each iteration reuses the standard forward pass to both evaluate the loss and extract dominant activation directions, without storing activation matrices across iterations.